# Testing the cosmological Euler equation: viscosity, equivalence principle, and gravity beyond general relativity

### v2

@@ VI Survey Specifications and Fiducials
In our forecast, each survey is analyzed by splitting the galaxy sample into two galaxy populations, namely bright and faint, determined by a redshift-dependent flux threshold. This threshold is chosen so that each population contains the same number of galaxies in each redshift bin. In our analysis, we assume a fiducial bias ratio of $b_{g,F}/b_{g,B}=0.5$ and choose the fiducial biases estimated by $b_{g,B}=1.4b_{g}(z)$ and $b_{g,F}=0.7b_{g}(z)$. This setup is closely aligned with the measurements reported in Bonvin:2023jjq. Moreover, a significant advantage of using the bias ratio is that it is observable, which can be derived by evaluating $\tilde{\beta}_{B}/\tilde{\beta}_{F}$, allowing us to directly obtain best-fit values and constraints once data become available.
For galaxy redshift uncertainties, they are modeled using a Gaussian velocity dispersion of $\sigma_{g}=4.24h^{-1}$ Mpc, following Howlett:2017asw. Magnification biases $s_{B,F}$ are computed based on the formalism in Bonvin:2023jjq and Maartens:2021dqy, while the evolution bias is set to zero at the fiducial, as it has been shown to produce subdominant effects from the observables in this work Bonvin:2023jjq; Castello:2023zjr. Throughout our analysis, we assume a fiducial cosmology with no violation of the EP and where $C_{\rm vis,0}=0$, i.e. $\tilde{E}_{P}=1$ at the fiducial. Moreover, the viscosity parameters are set to zero at the fiducial, namely $\mathcal{Z}=\mathcal{E}=0$.
For DESI BGS, the fiducial galaxy bias is set by the requirement $b_{g}^{\rm BGS}\,G(z)=1.34$, where $G(z)$ is the growth factor normalized to unity at the present epoch. The magnification biases for this sample are computed following Bonvin:2023jjq; Maartens:2021dqy. The adopted DESI BGS specifications are listed in Table 2.
Table 2: Survey Specifications for DESI BGS. Here and in the table below, the galaxy number density, $n_{g}$, is given in units of $10^{-3}\,(h/\mathrm{Mpc})^{3}$, assuming $h=0.67$. Survey volumes $V$ are expressed in $(\mathrm{Gpc}/h)^{3}$. We assume a fiducial bias ratio of $b_{g,B}/b_{g,F}=0.5$. The galaxy number density $n_{g}$ and the magnification bias $s_{B},s_{F}$, are obtained from Maartens:2021dqy. The observed absolute magnitude limit is chosen to be $m_{c}=19.5$. The evolution bias, $f_{evol}$, is set to zero.
@@
For Euclid, survey volume, fiducial galaxy biases, and the galaxy number density are taken from the Euclid paper Euclid:2019clj. Magnification biases are computed from the relations in Maartens:2021dqy, assuming a detector cut of $F_{c}=3\times 10^{-16}\,\mathrm{erg}\,\mathrm{cm}^{-2}\,\mathrm{s}^{-1}$. The Euclid specifications used in our forecasts are given in Table 3.
For SKA Phase 2, we adopt the aggressive survey configuration described in Quartin:2021dmr. Fiducial galaxy biases are derived using fitting functions for HI-selected galaxies Yahya:2014yva. Magnification biases are computed from the relations in Maartens:2021dqy, assuming a flux sensitivity limit of $5\,\mu\mathrm{Jy}$. The SKA Phase 2 specifications used in our forecasts are given in Table 4.
Table 3: Specifications for the Euclid spectroscopic survey, obtained from Euclid:2019clj and Maartens:2021dqy.
@@
Table 4: Specifications for SKA Phase 2, obtained from Bull:2015lja and Maartens:2021dqy.
@@
The magnification bias $s_{B,F}$ can be measured directly from the galaxy luminosity function (GLF). We follow the prescriptions in Maartens:2021dqy for all three surveys, and estimate the magnification bias for the B,F populations following the method of REF. We also estimate the uncertainty on $s_{B,F}$ by numerically propagating the errors of the GLF as provided in Pozzetti_2016 for Euclid and Loveday_2011 for DESI. We obtain the results of Table 5-7. For SKA2 we could not find the uncertainties of the GLF parameters, so we adopted the largest uncertainty of the DESI bins. Since in $\alpha_{F,B}$ we fixed $H$ to the fiducial, the corresponding errors are entirely due to $s_{F,B}$, and are estimated by standard error propagation. These uncertainties on $\alpha_{F,B}$, are then employed as priors in the Fisher analysis, since we treat $\alpha_{F,B}$ as free parameters here (instead of fixing them at the fiducials, as in Castello:2024lhl), which are fully degenerate with $\tilde{E}_{P}$.
@@ VII Results and Discussion
We present the forecasting results for the baseline bias configuration with $b_{g,F}/b_{g,B}=0.5$ in Tables 8, 9, and 10, corresponding to the DESI, Euclid, and SKA2 cases, respectively. All results displayed below are obtained by marginalizing over all the remaining parameters. We begin by commenting on the measurability of the viscosity parameter $C_{\rm vis,0}$. From these tables, we find that upcoming galaxy surveys will be capable of placing robust and remarkably precise constraints on the viscosity of dark matter. Namely, the present-day value of the dimensionless viscosity parameter $C_{\rm vis,0}$ can be constrained at the level of $\mathcal{O}(10^{-7})$, with the exact sensitivity depending on the specific survey configurations. We remark that our constraints are of a comparable order of magnitude to previous constraints (for a more detailed discussion on these constraints, see Appendix G), but with many fewer model assumptions.
@@
An additional parameter included in our Fisher analysis, which is not present in the previous work Castello:2024lhl, is the shot noise amplitude $n_{n}$. To allow for a more flexible and data-driven modeling of noise contributions that evolve with cosmic time, we treat $n_{n}$ as a free parameter in each redshift bin. As shown in Tables 810, we find that $n_{n}$ can be constrained to better than 1% across all redshift bins for both Euclid and SKA2. For DESI BGS, the constraints are slightly weaker, ranging from approximately 5% at low redshift to sub-percent precision at higher redshifts.
@@
To assess the impact of the bias ratio on the constraining power, we repeat the SKA2 forecast for three different bias configurations: the baseline choice $b_{g,F}/b_{g,B}=0.5$ (Table 10), a more asymmetric split with $b_{g,F}/b_{g,B}=0.2$ (Table 11), and a more symmetric case with $b_{g,F}/b_{g,B}=0.7$ (Table 12). We find that, consistent with expectations from our earlier analysis in Castello:2024lhl, larger differences in galaxy bias between the two tracers lead to tighter constraints on both $\tilde{E}_{P}$ and $C_{\rm vis,0}$. Specifically, the most asymmetric configuration ($b_{g,F}/b_{g,B}=0.2$) yields the tightest constraint on the viscosity parameter, with a $1\sigma$ uncertainty of $C_{\rm vis,0}<7.5\times 10^{-8}$, while the more symmetric case ($b_{g,F}/b_{g,B}=0.7$) leads to a weaker constraint of $\sim 1.7\times 10^{-7}$. These results highlight the importance of optimizing tracer selection to maximize sensitivity to relativistic effects and to break parameter degeneracies. We also find that the precision of $\tilde{E}_{P}$ is more sensitive to the bias ratio than that for $C_{\rm vis,0}$, reflecting its direct dependence on the dipole amplitude. Nevertheless, the viscosity parameter can be robustly constrained across a wide range of bias configurations.
@@
Lastly, we find that the constraints on $\tilde{E}_{P,z}$ are significantly weaker than those obtained in Castello:2024lhl for the viscous-free counterpart $E_{P}$. The key reason is that, in the present analysis, the parameters $\alpha_{B}$ and $\alpha_{F}$ are treated as free in each redshift bin. Since these parameters enter the dipole with the same structure as $\tilde{E}_{P,z}$, they are fully degenerate with it; thus, only external information can break the degeneracy. To remain conservative, we adopt broad Gaussian priors on $\alpha_{B}$ and $\alpha_{F}$, estimated from the uncertainties of the galaxy magnification biases (see, e.g., Table 7). Compared to fixed $\alpha$ values, this choice ensures robustness but inevitably weakens the constraining power on $\tilde{E}_{P,z}$. Future improvements in the determination of magnification biases will, therefore, be crucial for tightening the bounds on $\tilde{E}_{P,z}$.
@@ VIII Conclusion
Motivated by the question of whether viscosity would also interfere with the EP test proposed in Castello:2024lhl, we extended the previous model-independent analysis to incorporate viscosity. In addition to standard Newtonian terms, we took into account relativistic corrections in galaxy number counts, preserving all terms up to the order of $\lambda=\mathcal{H}/k$ in the galaxy power spectra. In this context, we identified the parameter $\tilde{E}_{P}$, which generalizes the original EP estimator $E_{P}$ in the presence of viscosity. Moreover, we showed that in the small-viscosity limit, the scale-independent component $\tilde{E}_{P,z}$ reduced to $E_{P}$ and can still serve as a smoking gun for EP violation.
@@
Looking ahead, the main conclusion of this paper points in two directions. First, the weak equivalence principle can still be tested using the null test on $\tilde{E}_{P,z}$ in the presence of dark matter viscosity, unless the true EP violation occurs only at the $10^{-6}$ level. In that latter case, one cannot disentangle EP violation from the effects of viscosity. Secondly, viscosity itself can be regarded as a promising quantity that can be robustly measured with the forthcoming Stage-IV surveys. Once real data become available, one can place constraints on it through an MCMC analysis. We leave these investigations for future work.
@@ Acknowledgments
LA acknowledges support by DFG under Germany’s Excellence Strategy EXC 2181/1 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster) and under Project 554679582 ”GeoGrav: Cosmological Geometry and Gravity with nonlinear physics”.
@@ A.3 Parametrization of the growth rate
In addition to the approximate solution for the growth rate given above, we found an analytic parametrization that extends the well-known relation $f=\Omega_{m}^{\gamma}$ (with $\gamma=0.545$) to the viscous case. We restrict ourselves to the case $\Theta=\Gamma=0$, $\eta=\mu_{\rm G}=1$. The contribution of viscosity to the background evolution was neglected.
@@ Appendix D Full expansion of the power spectra and parameters degeneracies
The coefficients $\mathcal{K}^{(BB)}_{a,b}$ and $\mathcal{K}^{(BF)}_{a,b}$, corresponding to the auto- and cross-power spectra $P_{\Delta_{B}\Delta_{B}}$ and $P_{\Delta_{B}\Delta_{F}}$, are listed in Tables 15 and 16, respectively. From these tables, we observe that, at first order in the expansion of $S_{g}^{2}$, the parameter $\sigma_{g}$ is fully degenerate with $\tilde{\beta}_{v}$ and $B_{v}$, and thus the $\sigma_{g}$ encoded in the exponential of $S_{g}$ is strongly degenerate with $b_{g}$. Therefore, in our analysis, we fix the $b_{g}$ encoded in $\tilde{\beta}_{v}$ to its fiducials. For a similar reason, we also fix $\mathcal{I}$ (as encoded in $B_{v},\tilde{\beta}_{B,v},\tilde{\beta}_{F,v}$, see Appendix A.2) to its fiducial values.
@@ Appendix F Viscous dark matter and coupled quintessence
$\qquad+\frac{3}{2}\Omega_{m}\Biggl(\mathcal{E}-\mathcal{Z}+\frac{\mathcal{Z}}{1-\mathcal{Z}}\Bigl((1+3\mathcal{Z})2Q^{2}+\mathcal{Z}Q\phi^{\prime}\Bigr)-\frac{1}{3}\frac{(\mathcal{Z}+\mathcal{E})\mathcal{E}}{1-\mathcal{Z}}\Bigl((1+3\mathcal{Z})Q\phi^{\prime}+\frac{1}{2}\mathcal{Z}\phi^{\prime 2}\Bigr)\Biggr)\Biggr]\theta_{m}$
@@
In terms of the notation introduced in Eq.s (25)-(27), we find
@@
$\quad+\frac{3}{2}\Omega_{m}\Biggl(\mathcal{E}-\mathcal{Z}+\frac{\mathcal{Z}}{1-\mathcal{Z}}\Bigl((1+3\mathcal{Z})2Q^{2}+\mathcal{Z}Q\phi^{\prime}\Bigr)-\frac{\mathcal{E}}{3}\frac{\mathcal{Z}+\mathcal{E}}{1-\mathcal{Z}}\Bigl((1+3\mathcal{Z})Q\phi^{\prime}+\frac{1}{2}\mathcal{Z}\phi^{\prime 2}\Bigr)\Biggr)$ (135)
@@
or writing in terms of $\Gamma,\Theta$ and $\tilde{\eta},\tilde{\mu}_{\rm G}$ (as defined in Appendix (I)),
@@
$\Theta$ $=(1+3\mathcal{Z})Q\phi^{\prime}+\frac{3}{2}\Omega_{m}\left[(\mathcal{Z}-\mathcal{E})\Bigl((1+3\mathcal{Z})2Q^{2}+\mathcal{Z}Q\phi^{\prime}\Bigr)-\frac{\mathcal{E}}{3}(\mathcal{Z}+\mathcal{E})\Bigl((1+3\mathcal{Z})Q\phi^{\prime}+\frac{1}{2}\mathcal{Z}\phi^{\prime 2}\Bigr)\right]$ (140)

### v3

@@ VI Survey Specifications and Fiducials
In our forecast, each survey is analyzed by splitting the galaxy sample into two galaxy populations, namely bright and faint, determined by a redshift-dependent flux threshold. This threshold is chosen so that each population contains the same number of galaxies in each redshift bin. In our analysis, we assume a fiducial bias ratio of $b_{g,F}/b_{g,B}=0.5$ and choose the fiducial biases estimated by $b_{g,B}=1.4b_{g}(z)$ and $b_{g,F}=0.7b_{g}(z)$. This setup is closely aligned with the measurements reported in Bonvin et al. (2023). Moreover, a significant advantage of using the bias ratio is that it is observable, which can be derived by evaluating $\tilde{\beta}_{B}/\tilde{\beta}_{F}$, allowing us to directly obtain best-fit values and constraints once data become available.
For galaxy redshift uncertainties, they are modeled using a Gaussian velocity dispersion of $\sigma_{g}=4.24h^{-1}$ Mpc, following Howlett et al. (2017). Magnification biases $s_{B,F}$ are computed based on the formalism in Bonvin et al. (2023) and Maartens et al. (2021), while the evolution bias is set to zero at the fiducial, as it has been shown to produce subdominant effects from the observables in this work Bonvin et al. (2023); Castello et al. (2024a). Throughout our analysis, we assume a fiducial cosmology with no violation of the EP and where $C_{\rm vis,0}=0$, i.e. $\tilde{E}_{P}=1$ at the fiducial. Moreover, the viscosity parameters are set to zero at the fiducial, namely $\mathcal{Z}=\mathcal{E}=0$.
For DESI BGS, the fiducial galaxy bias is set by the requirement $b_{g}^{\rm BGS}\,G(z)=1.34$, where $G(z)$ is the growth factor normalized to unity at the present epoch. The magnification biases for this sample are computed following Bonvin et al. (2023); Maartens et al. (2021). The adopted DESI BGS specifications are listed in Table 3.
Table 3: Survey Specifications for DESI BGS. Here and in the table below, the galaxy number density, $n_{g}$, is given in units of $10^{-3}\,(h/\mathrm{Mpc})^{3}$, assuming $h=0.67$. Survey volumes $V$ are expressed in $(\mathrm{Gpc}/h)^{3}$. We assume a fiducial bias ratio of $b_{g,B}/b_{g,F}=0.5$. The galaxy number density $n_{g}$ and the magnification bias $s_{B},s_{F}$, are obtained from Maartens et al. (2021). The observed absolute magnitude limit is chosen to be $m_{c}=19.5$. The evolution bias, $f_{evol}$, is set to zero.
@@
For Euclid, survey volume, fiducial galaxy biases, and the galaxy number density are taken from the Euclid paper Blanchard and others (2020). Magnification biases are computed from the relations in Maartens et al. (2021), assuming a detector cut of $F_{c}=3\times 10^{-16}\,\mathrm{erg}\,\mathrm{cm}^{-2}\,\mathrm{s}^{-1}$. The Euclid specifications used in our forecasts are given in Table 4.
For SKA Phase 2, we adopt the aggressive survey configuration described in Quartin et al. (2022). Fiducial galaxy biases are derived using fitting functions for HI-selected galaxies Yahya et al. (2015). Magnification biases are computed from the relations in Maartens et al. (2021), assuming a flux sensitivity limit of $5\,\mu\mathrm{Jy}$. The SKA Phase 2 specifications used in our forecasts are given in Table 5.
Table 4: Specifications for the Euclid spectroscopic survey, obtained from Blanchard and others (2020) and Maartens et al. (2021).
@@
Table 5: Specifications for SKA Phase 2, obtained from Bull (2016) and Maartens et al. (2021).
@@
The magnification bias $s_{B,F}$ can be measured directly from the galaxy luminosity function (GLF). We follow the prescriptions in Maartens et al. (2021) for all three surveys, and estimate the magnification bias for the B,F populations. We also estimate the uncertainty on $s_{B,F}$ by numerically propagating the errors of the GLF as provided in Pozzetti et al. (2016b) for Euclid and Loveday et al. (2011) for DESI. We obtain the results of Table 6-8. For SKA2 we could not find the uncertainties of the GLF parameters, so we adopted the largest uncertainty of the DESI bins. Since in $\alpha_{F,B}$ we fixed $H$ to the fiducial, the corresponding errors are entirely due to $s_{F,B}$, and are estimated by standard error propagation. These uncertainties on $\alpha_{F,B}$, are then employed as priors in the Fisher analysis, since we treat $\alpha_{F,B}$ as free parameters here (instead of fixing them at the fiducials, as in Castello et al. (2025)), which are fully degenerate with $\tilde{E}_{P}$.
@@ VII Results and Discussion
We present the forecasting results for the baseline bias configuration with $b_{g,F}/b_{g,B}=0.5$ in Tables 9, 10, and 11, corresponding to the DESI, Euclid, and SKA2 cases, respectively. All results displayed below are obtained by marginalizing over all the remaining parameters. We begin by commenting on the measurability of the viscosity parameter $C_{\rm vis,0}$. From these tables, we find that upcoming galaxy surveys will be capable of placing robust and remarkably precise constraints on the viscosity of dark matter. Namely, the present-day value of the dimensionless viscosity parameter $C_{\rm vis,0}$ can be constrained at the level of $\mathcal{O}(10^{-7})$, with the exact sensitivity depending on the specific survey configurations. We remark that our constraints are of a comparable order of magnitude to previous constraints (for a more detailed discussion on these constraints, see Appendix G), but with many fewer model assumptions.
@@
An additional parameter included in our Fisher analysis, which is not present in the previous work Castello et al. (2025), is the shot noise amplitude $n_{n}$. To allow for a more flexible and data-driven modeling of noise contributions that evolve with cosmic time, we treat $n_{n}$ as a free parameter in each redshift bin. This allows the noise contribution to deviate from the standard assumption that galaxies follow a Poisson sampling of the underlying matter distribution. This effective noise term can capture additional stochastic effects, such as observational systematics, selection effects (e.g., the Malmquist bias), or complex masking, which may vary across different redshifts and cannot be accurately inferred from galaxy number densities alone. As shown in Tables 911, we find that $n_{n}$ can be constrained to better than 1% across all redshift bins for both Euclid and SKA2. For DESI BGS, the constraints are slightly weaker, ranging from approximately 5% at low redshift to sub-percent precision at higher redshifts.
@@
To assess the impact of the bias ratio on the constraining power, we repeat the SKA2 forecast for three different bias configurations: the baseline choice $b_{g,F}/b_{g,B}=0.5$ (Table 11), a more asymmetric split with $b_{g,F}/b_{g,B}=0.2$ (Table 12), and a more symmetric case with $b_{g,F}/b_{g,B}=0.7$ (Table 13). We find that, consistent with expectations from our earlier analysis in Castello et al. (2025), larger differences in galaxy bias between the two tracers lead to tighter constraints on both $\tilde{E}_{P,z}$ and $C_{\rm vis,0}$. Specifically, the most asymmetric configuration ($b_{g,F}/b_{g,B}=0.2$) yields the tightest constraint on the viscosity parameter, with a $1\sigma$ uncertainty of $C_{\rm vis,0}<7.5\times 10^{-8}$, while the more symmetric case ($b_{g,F}/b_{g,B}=0.7$) leads to a weaker constraint of $\sim 1.7\times 10^{-7}$. These results highlight the importance of optimizing tracer selection to maximize sensitivity to relativistic effects and to break parameter degeneracies. We also find that the precision of $\tilde{E}_{P,z}$ is more sensitive to the bias ratio than that for $C_{\rm vis,0}$, reflecting its direct dependence on the dipole amplitude. Nevertheless, the viscosity parameter can be robustly constrained across a wide range of bias configurations.
@@
Lastly, we find that the constraints on $\tilde{E}_{P,z}$ are significantly weaker than those obtained in Castello et al. (2025) for the viscous-free counterpart $E_{P}$. The key reason is that, in the present analysis, the parameters $\alpha_{B}$ and $\alpha_{F}$ are treated as free in each redshift bin. Since these parameters enter the dipole with the same structure as $\tilde{E}_{P,z}$, they are fully degenerate with it; thus, only external information can break the degeneracy. To remain conservative, we adopt broad Gaussian priors on $\alpha_{B}$ and $\alpha_{F}$, estimated from the uncertainties of the galaxy magnification biases (see, e.g., Table 8). Compared to fixed $\alpha$ values, this choice ensures robustness but inevitably weakens the constraining power on $\tilde{E}_{P,z}$. Future improvements in the determination of magnification biases will, therefore, be crucial for tightening the bounds on $\tilde{E}_{P,z}$.
@@ VIII Conclusion
Motivated by the question of whether viscosity would also interfere with the EP test proposed in Castello et al. (2025), we extended the previous model-independent analysis to incorporate viscosity. In addition to standard Newtonian terms, we took into account relativistic corrections in galaxy number counts, preserving all terms up to the order of $\lambda=\mathcal{H}/k$ in the galaxy power spectra. In this context, we identified the parameter $\tilde{E}_{P}$, which generalizes the original EP estimator $E_{P}$ in the presence of viscosity. Moreover, we showed that in the small-viscosity limit, the scale-independent component $\tilde{E}_{P,z}$ reduced to $E_{P}$ and can still serve as a smoking gun for EP violation.
@@
Looking ahead, the main conclusion of this paper points in two directions. First, the weak equivalence principle can still be tested using the null test on $\tilde{E}_{P,z}$ in the presence of dark matter viscosity, unless the true EP violation occurs only at the $10^{-6}$ level. In that latter case, one cannot disentangle EP violation from the effects of viscosity. Secondly, viscosity itself can be regarded as a promising quantity that can be robustly measured with the forthcoming Stage-IV surveys. Once real data become available, one can place constraints on it through an MCMC analysis. It should be noted that the Fisher formalism is exact only if the likelihood is Gaussian and the observables depend linearly on all parameters. In practice, nonlinearities in the model mean the Fisher matrix provides only a local estimate of the covariance around the fiducial model. The true posteriors may have non-Gaussian tails or more complex features; thus, our Fisher forecast should be interpreted as a lower bound on the expected uncertainties from real data, representing the best-case statistical constraining power. We leave these investigations for future work.
@@ Acknowledgments
LA acknowledges support by DFG under Germany’s Excellence Strategy EXC 2181/1 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster) and under Project 554679582 ”GeoGrav: Cosmological Geometry and Gravity with nonlinear physics”. The Mathematica code we used to produce the Fisher matrix results is publicly available at 28.
@@ A.3 Parametrization of the growth rate
In addition to the approximate solution for the growth rate given above, we found an analytic parametrization that extends the well-known relation $f=\Omega_{m}^{\gamma}$ (with $\gamma=0.545$) to the viscous case. We restrict ourselves to the case $\Theta=\Gamma=0$, $\eta=\mu_{\rm G}=1$. Because the actual viscosity is so small (according to the constraints in Appendix G and the results of our analysis, the upper bound on $\mathcal{Z}$ is of order $10^{-5}$ to $10^{-6}$), the contribution of viscosity to the background evolution was neglected.
@@ Appendix D Full expansion of the power spectra and parameters degeneracies
The coefficients $\mathcal{K}^{(BB)}_{a,b}$ and $\mathcal{K}^{(BF)}_{a,b}$, corresponding to the auto- and cross-power spectra $P_{\Delta_{B}\Delta_{B}}$ and $P_{\Delta_{B}\Delta_{F}}$, are listed in Tables 16 and 17, respectively. From these tables, we observe that, at first order in the expansion of $S_{g}^{2}$, the parameter $\sigma_{g}$ is fully degenerate with $\tilde{\beta}_{v}$ and $B_{v}$, and thus the $\sigma_{g}$ encoded in the exponential of $S_{g}$ is strongly degenerate with $b_{g}$. Therefore, in our analysis, we fix the $b_{g}$ encoded in $\tilde{\beta}_{v}$ to its fiducials. For a similar reason, we also fix $\mathcal{I}$ (as encoded in $B_{v},\tilde{\beta}_{B,v},\tilde{\beta}_{F,v}$, see Appendix A.2) to its fiducial values.
@@ Appendix F Viscous dark matter and coupled quintessence
$\qquad+\frac{3}{2}\Omega_{m}\Biggl(\mathcal{E}-\mathcal{Z}+\frac{\mathcal{Z}}{1-\mathcal{Z}}\Bigl((1+3\mathcal{Z})2Q^{2}+\mathcal{Z}Q\phi^{\prime}\Bigr)-\frac{1}{3}\frac{(\mathcal{Z}+\mathcal{E})\mathcal{E}}{(1-\mathcal{Z})^{2}}\Bigl((1+3\mathcal{Z})Q\phi^{\prime}+\frac{1}{2}\mathcal{Z}\phi^{\prime 2}\Bigr)\Biggr)\Biggr]\theta_{m}$
@@
In terms of the notation introduced in Eqs. (25)-(27), we find
@@
$\quad+\frac{3}{2}\Omega_{m}\Biggl(\mathcal{E}-\mathcal{Z}+\frac{\mathcal{Z}}{1-\mathcal{Z}}\Bigl((1+3\mathcal{Z})2Q^{2}+\mathcal{Z}Q\phi^{\prime}\Bigr)-\frac{\mathcal{E}}{3}\frac{\mathcal{Z}+\mathcal{E}}{(1-\mathcal{Z})^{2}}\Bigl((1+3\mathcal{Z})Q\phi^{\prime}+\frac{1}{2}\mathcal{Z}\phi^{\prime 2}\Bigr)\Biggr)$ (135)
@@
or writing in terms of $\Gamma,\Theta$ and $\tilde{\eta},\tilde{\mu}_{\rm G}$ (as defined in Appendix I),
@@
$\Theta$ $=(1+3\mathcal{Z})Q\phi^{\prime}+\frac{3}{2}\Omega_{m}\left[(\mathcal{Z}-\mathcal{E})\Bigl((1+3\mathcal{Z})2Q^{2}+\mathcal{Z}Q\phi^{\prime}\Bigr)-\frac{\mathcal{E}}{3}\frac{\mathcal{Z}+\mathcal{E}}{1-\mathcal{Z}}\Bigl((1+3\mathcal{Z})Q\phi^{\prime}+\frac{1}{2}\mathcal{Z}\phi^{\prime 2}\Bigr)\right]$ (140)